OpenStudy (karinamtz0213):

Verify and show all steps needed please: 1.) cosx= 1/secx 2.) sin^2x-cos^2x=1 3.) cotx= sinx/cosx 4.) sinx= 1/ cscx 5.) 1+tan^2x=sec^2x 6.) 1/sinx=secx

OpenStudy (anonymous):

1, 4 and 6 are reciprocal identities. How does your teacher want you to verify them?

OpenStudy (karinamtz0213):

anyway possible as long as can be verifies correctly.

OpenStudy (anonymous):

Is it a matter of plugging in values? For example: sin 30° = 1/2 (y/r) csc 30° = 2 (r/y) so sin 30° = 1/csc 30°?

OpenStudy (karinamtz0213):

No. You have to make one of the sides equal the other side and no values are used.

OpenStudy (anonymous):

I'm rather new to trig functions, but one of our identities is sine is the reciprocal of 1/cosecant. I'm sorry ... I just don't know how you'd verify that. :(

OpenStudy (anonymous):

Oops... I meant that sine is the reciprocal of cosecant.